1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

The Banshee (An Approach)

Varulven Der Werwolf

The Werewolf

tradukita de Max Knight tradukita de Sven Collberg tradukita de Alexander Gross

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “In-

flect me, pray.”

En varulv lopp från

fru m.m. en vacker afton

samt begav sig till en bymagi-

sters grav och bad: “Var snäll

och konjugera.

Ein Werwolf eines

Nachts entwich von Weib und Kind

und sich begab an eines Dorfschul-

lehrers Grab und bat ihn: �Bitte,

beuge mich!“ A Werewolf, troubled by

his name, Left wife and brood one night and came To a hidden graveyard to

enlist The aid of a long-dead

philologist.

The village teacher

climbed up straight upon his grave stone

with its plate and to the apparition

said who meekly knelt be-

fore the dead: mig!” Byskolläraren

stod upp på bleckna-

mnsskyltens mässingknopp

med svar till den,

som satt på pass med tåligt korslagd

tass på tass.

Der Dorfschulmei-

ster stieg hinauf auf seines Blech-

schilds Messingknauf

und sprach zum Wolf, der seine Pfoten

geduldig kreuzte vor dem Toten:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit un-

steady Yawned twice, wheezed once, and then was ready.

“The banSHEE, in the

subject’s place; the banHERS, the

possessive case. The banHER, next, is

what they call objective case–and

that is all.”

“Jag var-ulv 1:a

pers. i sing., du, han, hon, den,

det var-ulv. Så vi voro-ulv, I voren.

På de voro-ulv kom in-

genting.”

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Geni-

tiv sodann, dem Werwolf, Dativ,

wie man’s nennt, den Werwolf, - damit

hat’s ein End!“ “Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

…  …  …  …

2/3

The banshee marveled

at the cases and writhed with plea-

sure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens

en enda pers. i tempus presens?”

Dem Werwolf schmeichelten die Fälle

er rollte seine Augenbälle.

�Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“ “I know that–I’m no men-

tal cripple– The future form and par-

ticiple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

The teacher, though,

admitted then that this was not wi-

thin his ken. “While bans are fre-

quent”, he advised, “A she cannot be plu-

rized.”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Hasbeen-wolf?’ Utterly wrong!

Such words are wounds beyond all suture– I’m sorry, but you have no

future.”

The banshee, rising

clammily, wailed: “What about

my family?” Then, being not a lear-

ned creature, said humbly “Thanks”

and left the teacher.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste exi-

stera.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

The Werewolf knew better–his sons still slept

At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .